OpenStudy (mtalhahassan2):

Hydrogen-3 has a half-life of 12.35 years. What mass of hydrogen-3 will remain form a 100.0 MG initial sample after 5.0 years? A) 7 MG B) 24 MG C) 76 MG D) 18 MG

OpenStudy (mtalhahassan2):

@wio

OpenStudy (mtalhahassan2):

@souvik

OpenStudy (mtalhahassan2):

Can someone please help me with it

OpenStudy (mtalhahassan2):

Plz

OpenStudy (astrophysics):

We will have to use the following \[\huge \frac{ m_f }{ m_i } = \left( \frac{ 1 }{ 2 } \right)^{\left( \frac{ t }{ T^{1/2} } \right)}\] where \[T^{1/2} = \text{half life}\] \[t =\text{ time}\] \[m_f = \text{final mass}\] \[m_i = \text{initial mass}\]

OpenStudy (astrophysics):

So let me ask you, what are we solving for here?

OpenStudy (mtalhahassan2):

Final mass

OpenStudy (astrophysics):

Right!

OpenStudy (astrophysics):

So our formula then becomes \[\huge m_f = m_i \left( \frac{ 1 }{ 2 } \right)^{\frac{ t }{ T^{\frac{ 1 }{ 2 }} }}\] right?

OpenStudy (astrophysics):

Now at this stage, you just need to plug in the values

OpenStudy (astrophysics):

and evaluate

OpenStudy (mtalhahassan2):

So what u get

OpenStudy (astrophysics):

I think you can at least do that much.

OpenStudy (mtalhahassan2):

Wait what we put for T 1/2

OpenStudy (astrophysics):

Don't let the 1/2 confuse you, I put it there to represent half life, that is 12.35 years

OpenStudy (mtalhahassan2):

Oh ok

OpenStudy (mtalhahassan2):

I am getting a very big number

OpenStudy (astrophysics):

\[\huge m_f = (100MG)(1/2)^{\frac{ 5.0 }{ 12.35 }}\]

OpenStudy (mtalhahassan2):

Know I get 76

OpenStudy (astrophysics):

Yup :)

OpenStudy (mtalhahassan2):

Thnx alot

OpenStudy (astrophysics):

Np